### 05 Apr 2026 v2

# High-resolution ultra-low-field MRI with SNRAware denoising

Abstract
Ultra-low-field (ULF, << 0.1 T) magnetic resonance imaging (MRI) systems offer advantages in cost, portability, and accessibility, but their current utility is still limited by low signal-to-noise ratio (SNR). Deep learning (DL)-based denoising has emerged as a potential strategy to mitigate this limitation. In this work, we present a systematic evaluation of a high-performance DL denoising model trained using the SNRAware framework and applied to 88 mT and 72 mT data. Using a series of controlled experiments, we assessed model performance as a function of spatial resolution, coil impedance matching, readout bandwidth, input noise level, $k$-space undersampling, anatomy, image contrast, and scanner platform, and compared against analytical denoising algorithms. The model consistently increased the effective SNR of ULF acquisitions, enabling images with nominal spatial resolutions comparable to those commonly used in clinical 3 T protocols. Residual analyses indicated that the model predominantly removed stochastic noise while preserving underlying signal structure. At the same time, the results highlight some constraints: denoising performance remains dependent on the starting SNR of the acquisition, and training-domain mismatch influences behavior under certain artifact conditions. These findings suggest that DL-based denoising can significantly expand the practical capabilities of ULF MRI, while emphasizing potential benefits from hardware-software co-optimization and the need for rigorous clinical validation to determine the diagnostic value of denoised images.
I Introduction
Noise in an image refers to random variations of brightness or color and is generally unwanted because it does not correlate with the spatial properties of the imaged object. In Magnetic Resonance Imaging (MRI), it often increases uncertainty in both diagnostic interpretation and quantitative analysis of medical images [1]. Noisy images result from reconstructing noisy data in the reciprocal spatial-frequency domain, $k$-space, where signals are electronic in origin. Random voltage fluctuations can stem from thermal processes in the sample or the resonant radio-frequency (RF) receive chain [2], or from electromagnetic interference (EMI) coupled into the system [3], all three leading to noisy $k$-space data. Importantly, their manifestation in the reconstructed images differs fundamentally from other confounding effects in MRI, such as aliasing, motion, eddy current distortions, off-resonance phenomena, or gradient non-linearities, which do not exhibit stochastic behavior [4].
The factors that determine the signal-to-noise (SNR) in MRI can be influenced at three different stages of the imaging process. Before acquisition, hardware design, grounding, coil engineering, shielding, impedance matching, and EMI suppression can reduce electronic noise, and in the best case the system can operate close to the so-called “thermal limit” [5]. During acquisition, imaging sequence timing, data averaging, bandwidth choices, and encoding strategies further influence the effective noise level [6]. Finally, after acquisition, noise can be reduced through dedicated denoising techniques applied either in image space or in $k$-space. The present work focuses on this latter stage, where post-processing strategies aim to recover the underlying signal (or reconstruction) from a noisy measurement.
In image space, classical (analytical) denoising methods consist of local and/or non-local filtering and regularization [7, 8]. In $k$-space, analytical denoising is often implemented through filtering or masking strategies, or through reconstruction approaches that produce denoised images from noisy data [9, 10, 11]. Deep learning (DL) methods have also been proposed for both image-space and reconstruction-integrated denoising, and have shown substantial improvements over traditional approaches in a wide range of MRI applications [12, 13]. Since the performance of DL-based denoisers depends strongly on their training strategy, SNRAware was recently introduced as a model-agnostic framework for training MRI denoising networks [14, 15]. Its key innovation is the incorporation of quantitative noise-distribution information obtained through SNR-unit reconstruction, together with g-factor based augmentation, to enhance the generalization when the testing data is out-of-distribution from the training data. SNRAware has demonstrated strong results at 1.5 T and 0.55 T across the 14 architectures evaluated in Ref. [14].
Ultra-low-field (ULF) MRI scanners can be made portable, affordable, and easy to deploy [16, 17, 18], although they are also inherently SNR-limited [19, 2]. Software denoising is therefore an attractive route to improve image quality [20], and both classical methods (e.g. Block-Matched 4-Dimensional filters, BM4D [21]) and DL approaches [22, 23, 24, 25] have been explored. Motivated by the performance of SNRAware at higher field strengths and its success to generalize to 0.55 T with all training data from 3 T, we hypothesized that it could also be valuable in the ULF regime, despite the substantial differences in the imaging hardware, statistical properties of noise, SNR levels, coil sensitivities, and contrast mechanisms encountered below 0.1 T.
In this paper, we conduct a systematic evaluation of a model trained with SNRAware to determine how it generalizes to the ULF setting. The denoiser network contains 200 million tunable parameters and was trained solely on 3 T data [15]. We examine its performance across a range of anatomies, contrasts, spatial resolutions, bandwidth conditions, noise environments, and sampling schemes, and compare it against both standard Fourier reconstructions and classical denoising. Our aim is to determine whether AI trained with high-field data can perform well at ULF-MRI, characterize the present operational limits, identify scenarios in which the model remains reliable, and highlight conditions under which its performance begins to degrade.
II Methods
II-A Model architecture and training
Figure 1: SNRAware training. Complex white noise is first sampled. Noise augmentation takes in the g-factor map to generate realistic, spatially variant, correlated noise. This is added to high-SNR data to create low-SNR images. The paired high-/low-SNR data is used to compute loss and optimized model parameters. In inference, complex, low-SNR images and g-factor maps are inputted into the trained model, which produces denoised complex images as final outputs. Further details can be found in Refs. [14, 15].
The denoising model used in this study was made available by Microsoft Research through a Technology Evaluation Program (TEP). This model is an imaging transformer and its architecture is featured with three types of attention modules purposefully designed to capture signal and noise characteristics across spatial and temporal dimensions in the imaging data. The inputs to the model are complex MR images and g-factor maps (trivial for single-coil receivers), and the output is the denoised complex image.
The $\approx 200$ million model weights were optimized within the SNRAware framework [14], a model-agnostic training methodology that incorporates quantitative information about noise statistics into the training process (Fig. 1). It uses SNR-unit reconstruction to normalize input images based on their underlying noise variance and incorporates a g-factor map as an additional channel to encode spatially variant noise amplification. The training set consisted of approximately three million cardiac cine images acquired at 3 T, augmented to simulate a wide range of SNR conditions by applying spatially varying and correlated noise sampled according to the measured g-factor distributions. Additional transforms included Partial Fourier (PF) emulation, $k$-space filtering, and coil-sensitivity-dependent intensity modulation. The model was trained on a cluster of 16 NVIDIA B200 GPUs with 183GB RAM over the course of one week, using a loss function that combined $\ell_{1}$ error in SNR-units with a structural similarity term. No additional fine-tuning or retraining was performed for the ULF data used in this study.
II-B MRI scanners
Figure 2: Portable Halbach scanners employed in this work. (a) 88 mT elliptical system for head and extremity imaging (NextMRI). (b) 72 mT circular system for extremity imaging (Physio I).
Most measurements were performed on a NextMRI system, a portable head and extremity MRI scanner (Fig. 2(a)) based on an 88 mT elliptical Halbach-array magnet [26]. The homogeneity of the main magnetic field is of $\approx 5,000$ ppm over a 20 cm diameter spherical volume, without cryogenic infrastructure and prior to active shimming. The scanner features a bore with an elliptical opening of 20 and 28 cm along the short and long axis, respectively. Gradient encoding is supplied by a set of three orthogonal coils driven by a commercial amplifier system, providing maximum amplitudes of approximately 45 mT/m and slew rates of 120 T/m/s. RF transmission (Tx) and reception (Rx) are performed with combined TxRx coils, and receive signals are routed through a low-noise amplifier chain optimized for operation at the 3.7 MHz Larmor frequency. System grounding, shielding, and cable routing were configured to minimize EMI during acquisition [5], although controlled EMI was introduced in specific experiments to assess model robustness.
The RF TxRx coil employed for all knee and phantom images was a solenoid with diameter and length both 150 mm, and quality factor $Q\approx 220$  [^1] The quality factor definition employed in this paper is $Q=\frac{\omega_{\text{c}}}{\Delta\omega_{\text{-3dB}}}$, with $\omega_{0}$ the center frequency and $\Delta\omega_{\text{-3dB}}$ the bandwidth between the -3 dB crossings of the magnitude plot of the parameter $S_{11}$ on either side of $\omega_{0}$.. For wrist imaging we employed an elliptical solenoid with inner/outer diameters of 100/140 mm, length 200 mm, and quality factor also $Q\approx 220$.
For a comparison across scanners, we employed a second portable MRI scanner, namely our Physio I system (Fig. 2(b)), a 72 mT circular Halbach system for extremity imaging [17].
II-C System control and data processing
In both scanners, pulse sequences were executed using MaRCoS [27, 28] with the MaRGE graphical interface [29]. Data were streamed through the Tyger reconstruction framework [30], which enabled remote processing of raw acquisitions and integration of SNRAware-based denoising into the standard reconstruction pipeline [31]. Processed images were returned to the scanner console through the same interface, allowing visualization in quasi-real time.
II-D Denoising performance experiments
TABLE I: Acquisition parameters for the RARE experiments used in phantom and in-vivo studies.
Experiment (Figure) FoV (mm3) # pixels PF (%) Res. (mm3) BW (kHz) Acq. time (ms) TR (ms) TE (ms) ETL TI (ms) $k$-space filling Scan time (min) Plane
Exp 1 (3a) $180\times 160\times 160$ $500\times 500\times 60$ 100 $0.3\times 0.3\times 3$ 42 12 200 20 5 center-out 20.0 AX
Exp 2: 2 mm (4) $180\times 160\times 160$ $90\times 80\times 80$ 100 $2\times 2\times 2$ 23 4 400 20 8 center-out 5.3 SAG
Exp 2: 1.5 mm (4) $180\times 160\times 160$ $120\times 112\times 112$ 100 $1.5\times 1.4\times 1.4$ 30 4 400 20 8 center-out 10.5 SAG
Exp 2: 1 mm (4) $180\times 160\times 160$ $180\times 160\times 160$ 75 $1\times 1\times 1$ 45 4 400 20 8 center-out 16 SAG
Exp 2: 0.2 mm (4) $150\times 150\times 150$ $750\times 750\times 36$ 100 $0.2\times 0.2\times 4.2$ 63 12 200 20 5 center-out 18 AX
Exp 3-4 (5a) $150\times 150\times 150$ $320\times 320\times 36$ 100 $0.5\times 0.5\times 4.2$ 160/27 2/12 200 20 5 center-out 7.7 AX
Exp 5 (5b) $150\times 150\times 150$ $320\times 320\times 36$ 100 $0.5\times 0.5\times 4.2$ 40 8 200 20 5 center-out 7.7 AX
Exp 6 (5c) $150\times 150\times 150$ $500\times 500\times 50$ 100/60 $0.3\times 0.3\times 3$ 42 12 200 20 5 center-out 16.7/10 AX
Exp 7: knee T1 (6a) Exp 8: (7a) $160\times 160\times 160$ $500\times 500\times 50$ 100 $0.3\times 0.3\times 3$ 42 12 200 20 5 center-out 16.7 AX/SAG/COR AX
Exp 7: knee STIR (6b) $\times 160\times 160 160×160$$\times 160\times 30$ 100 $1\times 1\times 5.3$ 40 4 1000 20 5 90 center-out 16 AX/SAG/COR
Exp 7: wrist T1 (6c) $160\times 80\times 120$ $400\times 300\times 26$ 100 $0.4\times 0.4\times 3.1$ 33 12 200 20 5 center-out 5.2 COR
Exp 7: wrist STIR (6c) Exp 8: (7b) $160\times 80\times 120$ $200\times 150\times 26$ 80 $0.8\times 0.8\times 3.1$ 17 12 1000 20 5 100/80 center-out 10.4 COR
Exp 9 (8) $150\times 150\times 150$ $320\times 320\times 36$ 100 $0.5\times 0.5\times 4.2$ 27 12 200 20 5 center-out 7.7 AX
Exp 10: T1 (9) $160\times 160\times 160$ $250\times 250\times 30$ 70 $0.6\times 0.6\times 5.3$ 42 6 200 20 5 center-out 3.5 SAG
Exp 10: STIR (9) $\times 160\times 160 160×160$$\times 160\times 30$ 60 $1\times 1\times 5.3$ 40 4 1000 20 5 90 center-out 9.6 SAG/COR/AX
Exp 10: T2 (9) $160\times 160\times 160$ $180\times 180\times 30$ 60 $0.9\times 0.9\times 5.3$ 30 6 1000 40 10 Linear 10.8 SAG/COR/AX
The main goal of this paper is to assess the impact of various operational aspects on the denoiser model performance. To this end, we acquired images in ten different experiments, which we describe below.
In all of them, two datasets were acquired during each measurement: the complex $k$-space signal and a noise-only acquisition with identical bandwidth and number of readout points. The noise measurement was performed at the beginning of the sequence while the magnetization reached its equilibrium value, and therefore did not increase the overall scan time. The standard deviation of the absolute value of the noise measurement ($\sigma_{\text{s}}$) was computed. The “raw” reconstructions ($\mathcal{I}_{\text{raw}}$) used throughout this work correspond to the inverse FFT ($\mathcal{F}^{-1}$) of the acquired $k$-space data ($s(k)$), normalized by $\sigma_{\text{s}}$ divided by the square root of the number of acquired voxels ($N_{\text{vox}}$, i.e. the total number of elements in the reconstruction, multiplied by the undersampling factor, see Exp. 6):
$\mathcal{I}_{\text{raw}}=\frac{\mathcal{F}^{-1}\{s(k)\}}{\sigma_{\text{s}}/\sqrt{N_{\text{vox}}}}=\frac{\mathcal{F}^{-1}\{s(k)\}}{\sigma_{\text{I}}}.$ (1)
This normalization converts the reconstructed images into SNR units [32]. These SNR-normalized images were then used as input to the denoising model, whose output ($\mathcal{I}_{\text{SNRA}}$) is expressed in the same scale.
To visually assess the characteristics of noise removal, we obtained difference (residual) images in some experiments by subtracting denoised magnitude images from the corresponding raw FFT magnitude reconstruction. In such cases, we provide SNR values in a region of interest (ROI), calculated as the mean voxel value divided by the standard deviation of those same voxel values. Likewise, we sometimes report noise estimations in image space by calculating the standard deviation of a ROI comprising exclusively background pixels. These can be used to confirm correct scaling to SNR units of $\mathcal{I}_{\text{raw}}$, where this measure should be $\sigma_{\text{I}}\approx 1$, and $\sigma_{\text{I,SNRA}}\ll 1$ for $\mathcal{I}_{\text{SNRA}}$.
All acquisitions used 3D-RARE pulse sequences. Sequence parameters and acquisition settings for all subsequent experiments are summarized in Table I.
II-D1 3 T clinical-resolution benchmark
Experiment 1 establishes a reference benchmark for model performance by T1-weighted (T1-w) knee imaging in an axial view at a spatial resolution commonly used in clinical 3 T protocols: 0.3 $\times$ 0.3 $\times$ 3 mm3.
II-D2 Model performance versus spatial resolution
Experiment 2 evaluated the performance of SNRAware as a function of spatial resolution. Sagittal knee acquisitions were performed with isotropic voxel sizes of 2 mm, 1.5 mm, and 1 mm using T1-w sequences. These resolutions progressively increase the SNR demands on the system while maintaining identical contrast preparation. To probe the model under a more extreme regime, an additional acquisition was performed on a structured ACR (American College of Radiology) phantom with an in-plane resolution of 0.2 $\times$ 0.2 mm2, requiring gradient amplitudes exceeding 40 mT/m.
II-D3 Model performance versus readout bandwidth
Experiment 3 evaluated the impact of readout bandwidth on denoising performance. At low field, the spectral response of the Rx chain can lead to spatial brightness modulation across the reconstructed image, particularly for short acquisition windows. Increasing the readout duration effectively narrows the acquisition bandwidth and reduces sensitivity to this spectral variation. For this experiment, we employed again the ACR phantom, under two readout conditions corresponding to acquisition windows of 2 ms (160 kHz) and 12 ms (27 kHz), respectively. All other sequence parameters were kept constant.
II-D4 Model performance versus impedance matching
Experiment 4 evaluated the sensitivity of the denoising model to variations in coil impedance matching. At low field, imperfect matching can also modify the spectral sensitivity profile of the Rx chain and reduce spatial brightness modulation in the reconstructed images. For this experiment, we employed again the ACR phantom, under two different matching conditions: an optimized configuration with reflected power below -20 dB, and a deliberately degraded configuration with reflected power of approximately -3 dB. All other acquisition parameters were kept constant.
II-D5 Model performance versus input noise level
Experiment 5 evaluated the robustness of the denoising model to variations in input noise amplitude. A baseline acquisition was first performed with the ACR phantom and the system was optimized to operate near thermal noise levels. An additional acquisition was then obtained after intentionally increasing environmental noise by placing a switched-mode power supply in close proximity to the receive coil, resulting in an approximately fivefold increase in the measured root-mean-square voltage noise. All other acquisition parameters were kept constant.
II-D6 Model performance versus undersampling
Experiment 6 evaluated the response of the denoising model to reduced $k$-space sampling. Undersampling can shorten acquisition times but introduces reconstruction-related artifacts and modifies the effective noise structure. Two sampling schemes were compared with the ACR phantom: a fully sampled acquisition and a Partial Fourier acquisition with 60 % filling in the second phase-encoding direction. All other sequence parameters were kept constant.
II-D7 Model performance versus anatomy, orientation, and contrast
Experiment 7 evaluated the generalization of the denoising model across different anatomical targets and imaging contrasts. To this end, we acquired T1-w, T2-w, and STIR (Short-Tau Inversion Recovery, often employed for fat signal suppression at low-field strengths [33]) images of a knee, as well as T1-w and STIR wrist images.
II-D8 Model performance versus BM4D
Experiment 8 compared SNRAware reconstructions against classical denoising. To this end, we applied BM4D filters to some FFT magnitude images and evaluated differences in SNR boosts, background noise, blurring, residuals, and $k$-space. The BM4D algorithm takes the standard deviation of the noise as an input parameter ($\sigma_{\text{BM4D}}$). This corresponds to $\sigma_{\text{I}}=1$ in $\mathcal{I}_{\text{raw}}$ if we consider the background noise, and a higher, brightness-dependent value for the different tissues. For this reason, we ran the BM4D algorithms with both $\sigma_{\text{BM4D}}=1$ and 2.
II-D9 Model performance versus scanner
Experiment 9 evaluated the sensitivity of the denoising model to differences between MRI systems. A T1-w acquisition of the ACR phantom was performed on two scanners operating at slightly different field strengths: the NextMRI system (88 mT) and the Physio I system (72 mT), both described in Sec. II-B.
II-D10 Clinical knee protocol at ULF
Experiment 10 evaluated the performance of the denoising model within a practical clinical imaging protocol based on a standard 3 T routine employed in a large radiology department. A multi-contrast knee protocol was designed to fit in an overall scan time of $\approx 40$ min while maintaining diagnostically relevant spatial resolution. The protocol included T1-w, T2-w, STIR acquisitions in multiple orientations. For this experiment, we employed the NextMRI scanner (88 mT). To meet the scan-time constraint, the spatial resolution of the acquisitions was slightly reduced with respect to Experiment 1 (see Table I).
III Results
Figure 3: Representative results from Experiment 1. Top: axial knee magnitude images reconstructed using FFT and SNRAware at 3 T clinical resolution (0.3 $\times$ 0.3 $\times$ 3 mm3). Middle: residual magnitude images relative to the raw FFT reconstruction. Bottom: magnitude of the central $k$-space slice for each reconstruction. The SNR (noise standard deviation) values displayed in the reconstructions were computed within the ROI indicated by the white (yellow) boxes.
Figure 3 shows axial knee images acquired at a spatial resolution of 0.3 $\times$ 0.3 $\times$ 3 mm3 and reconstructed using FFT and SNRAware, as described in Sec. II-D1. The top row presents the magnitude images ($\mathcal{I}_{\text{raw}}$, $\mathcal{I}_{\text{SNRA}}$), the middle row displays residual images computed by subtracting each denoised magnitude image from the corresponding raw FFT magnitude reconstruction, and the bottom row shows the magnitude of the central $k$-space slice associated with each reconstruction. The white boxes indicate the ROI where we computed the SNR for comparison, and the yellow boxes the ROI selected for background-noise characterization.
Figure 4: Representative results from Experiment 2. Sagittal knee images at isotropic resolutions of 2 mm, 1.5 mm, and 1 mm, and ACR phantom images at 0.2 $\times$ 0.2 mm2 in-plane resolution. Columns show raw FFT magnitude reconstructions, SNRAware-denoised images, and residual magnitude images relative to the raw FFT reconstruction. The SNR (noise standard deviation) values displayed in the reconstructions were computed within the ROI indicated by the white (yellow) boxes.
Figure 4 shows representative slices from sagittal knee and ACR phantom acquisitions at isotropic resolutions of 2 mm, 1.5 mm, and 1 mm, as well as an in-plane resolution of 0.2 $\times$ 0.2 mm2 for the phantom, as described in Sec. II-D2. The columns display the raw FFT magnitude reconstructions, the corresponding SNRAware-denoised images, and the residual magnitude images computed relative to the raw FFT reconstruction. Again, the white boxes indicate the ROI where we computed the SNR for comparison.
[…]
Figure 6: Representative results from Experiment 7. Raw FFT magnitude reconstructions and corresponding SNRAware-denoised images acquired under different anatomical and contrast conditions.
Figure 6 presents raw FFT magnitude reconstructions and the corresponding SNRAware-denoised images under the conditions described in Sec. II-D7 (Exp. 7). Panels (a) and (b) show, respectively, T1-w and STIR knee images along axial, sagittal, and coronal planes. Panel (c) displays a T1-w wrist acquisition (top), along with STIR images of the same wrist with two different inversion times: 100 ms (middle) and 80 ms (bottom).
Figure 7: Representative results from Experiment 8. Raw FFT magnitude reconstructions (left column) and the corresponding BM4D- and SNRAware-denoised images for the acquisitions described in Sec. II-D8. Panels (a) and (b) show knee axial T1-weighted and wrist STIR coronal images, respectively. The second and third columns display BM4D filtering with input parameters $\sigma_{\text{BM4D}}=1$ and 2, respectively, while the rightmost column shows the SNRAware reconstruction. For each case, the residual magnitude images and the central $k$-space slice are shown below the corresponding reconstruction. The SNR (noise standard deviation) values displayed in the reconstructions were computed within the ROI indicated by the white (yellow) boxes. The red arrows indicate regions where anatomical details appear different.
Figure 7 presents raw FFT magnitude reconstructions and the corresponding BM4D- and SNRAware-denoised images under the conditions described in Sec. II-D8 (Exp. 8). Here we have used the knee T1-w axial (a) and wrist 80 ms-STIR coronal (b) acquisitions from Experiment 7. The red arrows indicate regions where anatomical details differ in the close-ups. The second and third columns show the results of BM4D filtering with input parameters $\sigma_{\text{BM4D}}=1$ and 2, respectively. In all cases, we show the residuals and central $k$-space slices below the resulting images. The white boxes indicate the ROI where we computed the SNR for comparison, and the yellow boxes the ROI selected for background-noise characterization.
Figure 8: Representative results from Experiment 9. ACR phantom images acquired on two ULF-MRI systems: NextMRI (88 mT) and Physio I (72 mT). Columns show raw FFT magnitude reconstructions and SNRAware-denoised images. Zoomed-in panels display the high-resolution region of the ACR phantom. The SNR (noise standard deviation) values displayed in the reconstructions were computed within the ROI indicated by the red (yellow) boxes.
Figure 8 shows ACR phantom images acquired on two different MRI systems, as described in Sec. II-D9 (Exp. 9). The top row corresponds to the NextMRI scanner operating at 88 mT, and the bottom row to the Physio I scanner operating at 72 mT. Columns show raw FFT magnitude reconstructions and the corresponding SNRAware-denoised images, with a close-up view of the high-resolution region of the ACR phantom.
Figure 9: Representative results from Experiment 10. The raw images are shown in the top, with the corresponding denoised images below. This complete knee protocol lasts 40 min.
Figure 9 presents raw FFT magnitude reconstructions and the corresponding SNRAware-denoised images under the conditions described in Sec. II-D10 (Exp. 10). This 40 min protocol consists of: a T1-w sagittal image; STIR images with sagittal, coronal, and axial orientations; and a T2-w sagittal acquisition.
IV Discussion
IV-A 3 T clinical-resolution benchmark
The results of Experiment 1 demonstrate that SNRAware produces a substantial increase in effective SNR at ULF. In the white ROI shown in Fig. 3(a), the measured SNR increases from 3.2 in the raw FFT reconstruction to 38.8 after denoising ($12.1\times$ enhancement). This improvement enables acquisitions at 88 mT to reach nominal spatial resolutions commonly used in clinical 3 T protocols. It should be emphasized that matching the nominal voxel size does not imply equivalent image quality, since additional factors such as contrast, $B_{0}$ homogeneity, gradient linearity, eddy currents, and the intrinsically lower starting SNR at ULF still limit the attainable fidelity.
The noise standard deviation in the background ROI (yellow boxes) is $\approx 1.1$ in $\mathcal{I}_{\text{raw}}$ and $\approx 0.05$ in $\mathcal{I}_{\text{SNRA}}$. The former value is close enough to 1 to assume that scaling to SNR units as described in Eq. (1) was successful. Note that the spectral sensitivity of the coil modulates the background noise along the readout (horizontal) dimension (see Sec. IV-C), and we therefore expect the standard deviation of the noise to be slightly <1<1 towards the left and right edges of $\mathcal{I}_{\text{raw}}$, and slightly >1>1 at the center. Note also that the $\approx 12.1$ boost in the SNR of $\mathcal{I}_{\text{SNRA}}$ does not match the $\approx 22$ reduction in the noise standard deviation. This is consistent with $\mathcal{I}_{\text{raw}}$ being denoised according to the background standard deviation $\sigma_{\text{I}}$ (see Sec. II-D), which differs from the standard deviation of the pixel value at pixels where there is a non-zero signal, due to the Rician nature of noise in magnitude MR images [34].
SNRAware residuals feature quasi-pure noise, with the remaining vestigial structure exclusively hypointense and localized to cortical bone regions, where the underlying signal is intrinsically low and noise characteristics resemble those of the background. This behavior is again consistent with Rician noise.
The central $k$-space slices shown in Fig. 3 feature strongly suppressed peripheral noise with SNRAware relative to the raw reconstruction, with pronounced high-frequency structure emerging after the denoising.
IV-B Model performance versus spatial resolution
Experiment 2 reveals a marked dependence of the denoising performance on spatial resolution. Both the perceived image quality and the structure of the residual images in Fig. 4 appear to improve as the voxel size decreases and approaches the resolution regime typical of clinical MRI. One possible explanation is that the SNRAware-trained model learned exclusively from high-resolution clinical datasets, and therefore operates closer to its native domain at smaller voxel sizes.
[…]
The phantom experiment provides a useful test of the qualitative behavior of the denoising model in ULF conditions. Because the training data consisted of 3 T clinical images, the model was never exposed to artifacts commonly observed at ULF. These include zipper artifacts produced by EMI with discrete spectral signatures, as well as brightness modulation along the readout direction arising from the frequency-dependent sensitivity of the RF Rx chain in high-bandwidth acquisitions (see Sec. IV-C). Generative AI models may suppress or hallucinate such features when they conflict with learned priors, effectively performing operations beyond denoising [35]. In contrast, the SNRAware model preserves both the zipper artifact and the brightness modulation in the reconstructed images. The corresponding residual images contain predominantly noise, with different magnitudes inside and outside the phantom, consistent with the expected behavior of a model performing noise suppression without altering the underlying signal structure.
IV-C Model performance versus readout bandwidth
The results of Experiment 3 are shown in Fig. 5(a). High-bandwidth (short-readout) acquisitions lead to: i) higher background noise in $\mathcal{I}_{\text{raw}}$, which grows as the square root of the readout bandwidth; ii) a more distinct intensity modulation along the readout direction due the spectral sensitivity of the RF receive chain; iii) the appearance of a zipper artifact due to pick up of a narrow-band EMI; and iv) suppressed distortions with respect to the lower bandwidth acquisition. The effect of the denoiser is notorious in both cases, but the high-resolution region of the ACR phantom is only resolved with the longer readouts.
IV-D Model performance versus impedance matching
The results of Experiment 4 are also shown in Fig. 5(a). Comparing the two columns corresponding to the impedance matching conditions reveals that the acquisition with reflected power below 20-20 dB produces higher image quality than the deliberately degraded matching condition at approximately 3-3 dB. In the zoomed views of the high-resolution region of the phantom, the structures on the left-hand panels (20-20 dB matching) appear more clearly resolved than those on the right-hand panels (3-3 dB matching) for the longer acquisition, whereas both are heavily blurred for the shorter 2 ms readouts, as the initial SNR is substantially lower.
When the bandwidth is high (short acquisition window), degrading the impedance matching can improve the visual uniformity of the reconstructed image. The intensity modulation becomes less pronounced when the spectral response of the coil is flattened.
Although operating closer to optimal impedance matching is mostly advantageous, this difference in performance becomes relatively small after denoising. This observation may have practical implications for scanner operation, as relaxing the requirement for precise tuning and matching before every acquisition could simplify workflows and reduce the need for repeated adjustments between patients or even between individual scans.
IV-E Model performance versus input noise level
[…]
IV-F Model performance versus undersampling
The effect of $k$-space undersampling (Exp. 6) is shown in Fig. 5(c). The SNRAware model produces visually coherent images for both the fully sampled acquisition and the PF acquisition with a filling factor of 0.6. Nevertheless, the fully sampled reconstruction yields overall better image quality, consistent with the higher information content of the acquired $k$-space data.
In the zoomed panels of the high-resolution region of the phantom, the grid-like structures in the middle and right sections of the phantom appear slightly more visible in the PF reconstruction than in the fully sampled case. While this behavior may occur occasionally due to the particular sampling pattern and noise realization, repeated observations suggest that fully sampled acquisitions generally lead to better-resolved structures. However incidental, these results reveal an informative aspect of the model behavior. In the fully sampled reconstruction, SNRAware largely suppresses the middle and right grid structures, treating them as noise-like features, similar to what was observed in the elevated-noise experiment in Fig. 5(b). In contrast, in the PF reconstruction the corresponding regions appear as unresolved blurred patterns that remain distinguishable from the surrounding background noise. The model has preserved low-confidence structures rather than removing them entirely, but there seems to be more information content in the raw close-up than in the denoised image.
IV-G Model performance versus anatomy, orientation, and contrast
The results of Experiment 7 demonstrate that SNRAware provides substantial SNR improvements across a broad range of anatomical targets, imaging orientations, and contrasts. As shown in Fig. 6, the denoised reconstructions exhibit visibly higher SNR than the corresponding raw FFT images for both knee and wrist acquisitions, and across T1-w, T2-w, and STIR contrasts. Similar behavior has also been reported in ULF brain imaging [31]. The improvement is consistent across axial, sagittal, and coronal orientations, indicating that the model generalizes well to variations in anatomical structure and image geometry. The amount of detail recovered in the STIR image ($\text{TI}=80$ ms) denoised with the model is particularly remarkable. The BM4D denoised images also recover some of the bone structure, but to a lesser extent.
IV-H Model performance versus BM4D
From the results of Experiment 8 it is apparent that classical BM4D filtering also provides a strong improvement in terms of SNR in the selected ROIs, with boosts ranging from 3.1 to 26.9 ($8.7\times$) for T1-w, to 2.5 to 26.2 ($10.5\times$) for STIR (see Fig. 7). These are to be compared with SNR boosts of $9.4\times$ and $15.6\times$ in the corresponding DL-denoised reconstructions. However, this is for the case where $\sigma_{\text{BM4D}}=2$, which leads to significant blurring and loss of anatomical detail (see red arrows). These effects are less evident with $\sigma_{\text{BM4D}}=1$, but the denoising performance is significantly lower, with SNR increases of only $1.9\times$ and $5.9\times$ for the T1-w and STIR images, respectively. These results show strong advantages of SNRAware with respec to BM4D in low-SNR imaging regimes.
The central $k$-space slice also points at better performance for SNRAware than the BM4D images, with much more structured content in both the T1-w and the STIR acquisition, with the exception of the $\sigma_{\text{BM4D}}=2$ case with STIR, where many features are also visible across the whole $k$-space slice.
IV-I Model performance versus scanner
The results of Experiment 9 compare ACR phantom acquisitions obtained with the NextMRI system (88 mT) and the Physio I scanner (72 mT). In both cases, the systems were operated after optimizing the hardware configuration to reach noise levels compatible with the thermal limit, so the observed difference is consistent with the modest increase in field strength.
As shown in Fig. 8, the images acquired with the NextMRI scanner exhibit slightly higher quality than those obtained with Physio I. This is visible in the high-resolution region of the phantom, even if the SNR in the ROI is higher for the lower field strength. The effect is more pronounced in the raw FFT reconstructions than in the denoised images. After applying SNRAware, the visual differences between the two systems become less pronounced, indicating that the denoising process partially compensates for the SNR advantage associated with the higher magnetic field.
The SNR boosts at 88 and 72 mT are $4.3\times$ and $14.1\times$, respectively.
As a final note, the NextMRI image appears more distorted than with Physio I due to the higher field inhomogeneity, as we have not applied any distortion correction measure for these experiments [36, 31].
IV-J Clinical knee protocol at ULF
Figure 9 (Exp. 10) illustrates a practical imaging protocol optimized to fit within a clinically reasonable scan time of approximately 40 min. This protocol includes five sequences for T1-w, T2-w, and STIR acquisitions and is the basis of an ongoing clinical study of 134 subjects with knee pathologies. To meet the scan-time constraint, the spatial resolution of the acquisitions was slightly reduced with respect to Experiment 1 (see Table I).
V Conclusions & Outlook
This work presented a systematic evaluation of DL-based denoising for ULF-MRI using a model trained with the SNRAware framework. Across a series of controlled experiments, the model demonstrated the ability to substantially increase the effective SNR of ULF acquisitions. In particular, Experiment 1 showed that images acquired at 88 mT can reach nominal spatial resolutions comparable to those used in clinical 3 T protocols, albeit without implying equivalence in overall image fidelity. The experiments further revealed that the denoising model performs consistently across a wide range of acquisition conditions, including different spatial resolutions, anatomical targets, contrasts, orientations, and scanner platforms.
A central observation emerging from this study is the ability of the model to remove noise while preserving the underlying signal structure. Residual images obtained by subtracting the denoised reconstructions from the raw images typically exhibit noise-like patterns with minimal structured content, indicating that the model acts predominantly as a noise suppressor rather than performing implicit image restoration or hallucination. Given the fundamentally stochastic nature of noise, the ability of an algorithm to identify and remove noise contributions in this manner may appear counterintuitive. The results presented here nevertheless demonstrate that DL-based models trained with appropriate noise statistics can perform this task with remarkable effectiveness. This performance likely stems from the SNRAware training strategy, which leverages temporal redundancy in cardiac cine data to distinguish signal components that are consistent across frames from noise that is uncorrelated. When applied to 3D ULF acquisitions, a similar separation appears to take place across slices, where spatially coherent structures are preserved while uncorrelated components are efficiently suppressed, even for extremely low-SNR acquisitions (see Fig. 7(b)).
At the same time, the experiments highlight several limitations. The quality of the denoised output remains dependent on the starting SNR of the acquisition, as illustrated by the elevated-noise experiment (Fig. 5(b)). Likewise, the performance of the model appears to improve as the spatial resolution approaches the regime of the training data, suggesting that training-domain mismatch plays a role in determining the model’s behavior (Fig. 4). Hardware factors such as impedance matching, bandwidth selection, and environmental noise also influence the final reconstruction quality, emphasizing that denoising performance ultimately reflects the combined properties of both the acquisition system and the post-processing algorithm.
Despite these limitations, the SNR gains observed in this work are sufficiently large to potentially alter the practical capabilities of ULF-MRI systems. Poor SNR has long been considered one of the primary limitations of low-field imaging, often competing with $B_{0}$ inhomogeneity as the dominant constraint on image quality [2, 20]. The results presented here indicate that DL-based denoising can mitigate this limitation to a considerable extent, enabling substantially higher spatial resolution imaging than previously achievable at these field strengths. It is therefore plausible that the principal bottleneck in ULF-MRI shifts from SNR to acquisition time, emphasizing the need for improved sampling efficiency and acceleration strategies.
Further progress will benefit from training models on datasets that explicitly include artifacts commonly observed in ULF systems, such as EMI patterns and receive-chain spectral modulation. Incorporating these effects into the training domain may allow DL-based denoisers to address both stochastic noise and structured acquisition artifacts, further expanding the applicability of these methods in ULF-MRI.
At the same time, the clinical implications of DL-based denoising must be rigorously assessed. The present study has taken a broad and systematic approach to characterizing the behavior of the model across a wide range of acquisition conditions, but the analysis remains largely qualitative and focused on technical image properties. The ultimate test of these methods will therefore be their performance in clinical practice. This will require acquiring images from patient cohorts and evaluating them through blinded analysis by specialized radiologists in order to determine the true diagnostic value of DL-denoised ULF images. A first such study is presently under way and may establish whether the substantial SNR gains demonstrated here translate into meaningful improvements in clinical decision-making.
Contributions
See Table II.
TABLE II: Author contributions. An “x” indicates participation in the corresponding task.
Task TGN HX JMA EGC JC FG MAN JS LVC MH JA
Scanner preparation x x x x x x
SNRAware/Tyger prep. x x x
Data acq. x x x x x x x
Data processing x x x
Proj. conception x x x x
Proj. manage. x x x
Figure prod. x x x x
Paper writing x x x
Paper revision x x x x x x x x x x x
Acknowledgment
We thank Luis Martí-Bonmatí and the team at Instituto de Investigación Sanitaria at La Fe Hospital for sharing the clinical knee protocol at 3 T we based on for Experiment 10.
Funding
This work was supported by Ministerio de Ciencia e Innovación (PID2022-142719OB-C22) and the European Innovation Council (NextMRI 101136407). JC acknowledges funding from the Spanish Ministry of Science, Innovation and Universities through a Formación del Profesorado Universitario grant (FPU23/01559).
Code availability
All material related to the denoising network (SNRAware) is available at https://github.com/microsoft/SNRAware/.
Conflict of interest
TGN consults for PhysioMRI Tech. JMA, FG, and JA are co-founders of PhysioMRI Tech.